Jdiaz45:

Which equation is y=-3x^2-12x-2 rewritten in vertex form

dude:

Make sure to include the available choices

Tranquility:

Hello! Welcome to QuestionCove! This is the vertex form y = a(x - h)^2+ k where (h, k) is the vertex To find the x-coordinate of the vertex it's -b/2a when the equation is in the form ax^2 + bx + c You can then plug in the x-coordinate of the vertex into the equation to find the y-coordinate of the vertex. You then plug in those two values into the vertex form of the equation

Ferredoxin4:

Good start, but find the vertex to plug into the vertex form does not guarantee that a=1. This is why you must apply complete the square to convert standard form to vertex form: \[-3x^2-12x-2\] \[-3(x^2+4x+c)=2+(-3c)\] Apply what you know about perfect square trinomials and c=4 since (b/2)^2 = 2^2. \[-3(x^2 + 4x + 4)=2-12\] \[-3(x+2)^2 = -10\] \[y=-3(x+2)^2 +10\]

Tranquility:

Good point, completely slipped my mind. Thank you for the correction